OpenStudy (anonymous):

The point (2,−11) is reflected in the x-axis and then reflected in the line y=x to get the point (a,b). What is the value of a+b?

OpenStudy (anonymous):

13 is the answer. You want the detail?

OpenStudy (unklerhaukus):

who cares about the answer , all i want is the detail

OpenStudy (anonymous):

yes, detail about your answer give me more confidence

OpenStudy (anonymous):

When you reflect a point in the x- axis the x co-ordinate stays the same and the y co-ordinate becomes negative:  So (2, -11) will become (2, 11). Now there are 2 ways to proceed. You can either say that when you reflect a point in y=x you swap the co-ordinates, so the answer will be (11, 2), or you can go for a more complete answer: When we reflect (2, 11) in the line y=x to get (a, b), the mid point of (a, b) and (2, 11) will lie on the line y=x. So the point: \[(\frac{a+2}{2}, \frac{b+11}{2})\] Will lie on the line y=x. Hence we have: \[\frac{a+2}{2} = \frac{b+11}{2}\] and so \[a+2=b+11\] So: \[a=b+9 \] Let's call this equation 1. Now, another piece of information we can gain: the slop between the points (a, b) and (2, 11) will be -1. So use the gradient formula: \[\frac{b-11}{a-2}=-1\] \[b-11=2-a\] \[b=13-a\] Substituting this into equation 1 we have: \[a=(13-a)+9\] \[2a=22\] \[a=11\] Now substituting a=11 into b=13-a gives b=2. So (a, b)=(11, 2) and a+b=13.

OpenStudy (anonymous):

@irudayadhason is there anything I can explain more for you?

OpenStudy (anonymous):

no.i am going to type a new question please answer that, i got it for this question

OpenStudy (anonymous):

please don't reply for this question, it's taking a long time i want to type my next question.

OpenStudy (anonymous):

Okidoke

OpenStudy (anonymous):

please demitris don't reply for this question, it's taking a long time i want to type my next question